# MIMIR: Masked Image Modeling for Mutual Information-based Adversarial Robustness

### v3

@@ 3.3 Design Details
Autoencoder. MIMIR follows the general design of MAE (He et al. 2022), consisting of an encoder $f_{e}$ and a decoder $f_{d}$. As with other autoencoders, the encoder extracts discriminative features $z$ from inputs $x$. The decoder reconstructs original inputs according to the discriminative features. Following ViT (Dosovitskiy et al. 2021), the input $x$ is separated into non-overlapping image patches. We randomly mask out a part of the patches and use the remaining patches as inputs for the following process in the encoder. This random masking process uses uniform distribution to prevent potential sampling bias, such as all foreground information being masked, as it becomes infeasible to find the reconstruction target. Thus, we aim to keep a part of the foreground as a hint for reconstruction. The information of masked content is recorded as mask tokens $m$, not used by the encoder but reserved for later use by the decoder. Each token is a learned vector indicating the presence of a masked patch to be predicted. The mask token is shared by all inputs of the decoder. Like unmasked patches, mask tokens are also assigned corresponding positional embeddings to be in the correct location in the reconstructed image. We emphasize that mask tokens are not used in the encoder part.
@@
To train a ViT, we use the same transformer blocks as ViT to build the encoder. The encoder only processes the visible patches, making training much more efficient. When converting to other architectures, such as ConViT (D’Ascoli et al. 2021), we use corresponding transformer blocks to build the encoder. The decoder accepts the encoded visible image patches and mask tokens as inputs. The decoder is built using the same transformer blocks as the encoder instead of using ViT (Dosovitskiy et al. 2021) transformer blocks for all. Then, the decoder is followed by a fully connected layer, which outputs the same number of patches as the original image.
@@
Adversarial Pre-training Target. The training target is to extract discriminative features from visible image patches by the encoder and then reconstruct the invisible patches by the decoder. Therefore, we need a differentiable measurement to quantify the distance between the original image and the reconstructed results. Following the original MAE (He et al. 2022), this distance is measured by Mean Squared Error (MSE). To create a more difficult reconstruction task, we apply adversarial perturbations $\delta$ on the inputs of the encoder. Thus, the adversarial perturbations are also masked along with the image upon input into the encoder. The decoder reconstructs the original natural inputs by using the latent features $z$ extracted from adversarial examples. The outputs of decoder $x_{re}$ and $x$ are used to calculate the MSE loss, which is further used to optimize the model. Note that the reconstruction differs from the original MAE; we do not use the encoder inputs as reconstruction targets. Formally, the pre-training process (described in Algorithm 1) can be written as follows:
@@
MI as Penalty. Inspired by IB, we show in Section 3.4 that MI between latent representation and adversarial examples decreases as the accuracy on adversarial examples, i.e., $I(x+\delta,z)$ is decreasing while training. Motivated by this finding, we directly use $I(x+\delta,z)$ as a penalty in our final loss function:
@@
where $\lambda$ is a regularizer for the MI penalty. We use $I(x+\delta,z)$ instead of $I(x,z)$ as a penalty. This is because $x\to x+\delta\to z$ follows the Markov chain since $z$ is extracted from $x+\delta$. According to Data Processing Inequality (DPI) (Beaudry and Renner 2011), $I(x,z)\leq I(x+\delta,z)$. $I(x+\delta,z)$ is closer to $z$ on the Markov chain.
Generating Adversarial Examples. To conduct the adversarial pre-training, we need an attack that finds proper adversarial perturbations $\delta$. As the autoencoder does not provide classification outputs, it is not possible to directly use existing adversarial attacks, such as PGD (Madry et al. 2018). Nevertheless, it is feasible to design a new algorithm to find $\delta$ by maximizing $loss_{\text{mse}}$ in Eq. (4). As the feature $z$ is extracted from only visible image patches, we only attack the visible patches. We do not add any perturbations to mask tokens since the outputs of the autoencoder are only impacted by visible patches. Then, the adversarial pre-training learning objective can be written as:
@@
Figure 2: The example plots for the lower and upper bounds on the MI in Propositions 3.2 and 3.3. The entropy ($H(\cdot)$) is chosen uniformly at random from a set of 10 classes. The lower bound reaches its minimum at $p_{e}=0.9$.
3.4 Theoretical Justification
Next, we provide theoretical justification showing that the MI between the adversarial example and its latent representation, i.e., $I(x+\delta,z)$, should be constrained. Let $F$ denote any classifier trained on natural samples with desirable prediction accuracy, which may suffer from adversarial attacks. We begin our analysis by first presenting Lemma 3.1.
Lemma 3.1.
@@ 3.4 Theoretical Justification
Now, we define $p_{e}$ as the probability that the predicted label of $x+\delta$ by $F$ is not equal to that of $x_{re}$, i.e., $p_{e}=\mathbb{P}(F(x+\delta)\neq F(x_{re}))$. Intuitively, our autoencoder is trained to recover only natural sample $x$ without any interference from $\delta$. Hence, a relatively large value of $p_{e}$ is expected. In the following, we establish the connection between $p_{e}$ and $I(x+\delta,z)$ with both lower and upper bounds, showing that minimizing $I(x+\delta,z)$ also encourages a large value of $p_{e}$.
[…]
Architecture Params (M) FT FT Epoch Adv. Steps Source Natural Robust
ResNet-50 25.0 PGD 100 1 (Bai et al. 2021) 67.44 35.54
DeiT-S 22.1 PGD 100 1 (Bai et al. 2021) 66.62 36.56
DeiT-S 22.1 PGD 110 1 (Debenedetti, Sehwag, and Mittal 2023) 66.80 37.90
ViT-S 22.1 PGD 120 3 (Mao et al. 2022) 66.43 39.20
ViT-S 22.1 PGD 300 3 (Liu et al. 2024) 70.7 43.7
RobArch-S 26.1 PGD 110 3 (Peng et al. 2023) 70.17 44.14
ViT-S 22.1 APGD 300 2 (Singh, Croce, and Hein 2023) 69.22 44.04
ViT-S 22.1 APGD 300 2 Ours 71.00 46.10
ViT-B 86.6 ARD+PRM 10 5 (Mo et al. 2022) 69.10 34.62
ViT-B 86.6 PGD 120 3 (Mao et al. 2022) 70.64 43.04
RobArch-L 104 PGD 100 3 (Peng et al. 2023) 73.44 48.94
ViT-B 86.6 PGD 300 3 (Liu et al. 2024) 74.7 49.7
[…]
ViT-B 86.6 APGD 300 2 (Singh, Croce, and Hein 2023) 74.10 50.30
ViT-B 86.6 APGD 300 2 Ours 76.32 54.28
Proposition 3.2.
@@
By applying Fano’s inequality (Fano 1949; Ocal et al. 2019), we obtain:
@@
The last line of Eq. (14) is by Lemma 3.1. ∎
Therefore, we obtain a lower bound of $I(x+\delta,z)$. If we use CIFAR-10 ($|F(x+\delta)|=10$) and assume the predicted labels $F(x+\delta)$ follow a uniform distribution, we can visualize the lower bound as a function of $p_{e}$ as shown in Figure 2, from which we observe an obvious monotonic inverse relationship between $I(x+\delta)$ and $p_{e}$ in the range $p_{e}\in[0,0.9]$. In fact, we can also obtain an upper bound, under the assumption that $I(F(x+\delta),x_{re})\approx I(x+\delta,z)$, i.e., there is no information loss in the two Markov chains in Eq. (8).
Proposition 3.3.
@@
By the Hellman-Raviv inequality (Hellman and Raviv 1970; Brown 2009), we have:
@@
Similar to the lower bound, the upper bound also indicates $I(x+\delta,z)$ is inversely proportional to $p_{e}$ as shown in Figure 2. In fact, apart from the above-mentioned lower and upper bounds, there exists an alternative and intuitive way to understand the mechanism of minimizing $I(x+\delta,z)$. For simplicity, let us assume the natural data $x$ and adversarial perturbations $\delta$ are independent [^3] This assumption is mild for certain scenarios, such as when considering universal or image-agnostic perturbations (Moosavi-Dezfooli et al. 2017)., then:
@@
According to (Vincent et al. 2008), minimizing the expected reconstruction error between natural sample $x$ and corrupted input $x+\delta$ amounts to maximizing a lower bound of the mutual information $I(x,z)$, even though $z$ is a function of the corrupted input. Therefore, by minimizing $I(x+\delta,z)$, the network is forced to minimize $I(\delta,z)$ (since $I(x,z)$ is maximized). In other words, only the adversarial information about $\delta$ has been removed from $z$ when minimizing $I(x+\delta,z)$. This also explains the robustness of $z$.
4 Experiments
4.1 Experimental Setup
@@
We evaluate MIMIR on three datasets: ImageNet-1K (Deng et al. 2009), Tiny-ImageNet (Le and Yang 2015), and CIFAR-10 (Krizhevsky, Hinton et al. 2009). Details of datasets are provided in Appendix A. Hyperparameters of the decoder are included in Appendix B.
@@
Training Setup. We train models from scratch for all experiments. Following MAE (He et al. 2022), we do pre-training by MIMIR for 800 epochs. The number of warmup epochs is 40 for pre-training. We use AdamW (Loshchilov and Hutter 2019) as an optimizer for both pre-training and fine-tuning. The cosine decay as the learning rate scheduler is applied. At pre-training, MIMIR uses the 1-step PGD to generate adversarial examples for all three datasets. The perturbation budget is $\epsilon=8,\alpha=10$. At fine-tuning stage, we use the 10-step PGD AT with perturbation bound $\epsilon=8,\alpha=2$ for CIFAR-10 and Tiny-ImageNet. For ImageNet-1K, we use 1-step PGD with random initialization for better efficiency. The perturbation bound is $\epsilon=4$ for ImageNet-1K. We also use a 2-step APGD with a longer fine-tuning scheduler to compare with SOTA works in Table 1. Details on training hyperparameters are provided in Appendix C.
Evaluation Metrics. We use natural accuracy and robust accuracy as evaluation metrics. Natural accuracy refers to the accuracy of natural and unmodified inputs. The robust accuracy measures the accuracy under the AutoAttack (Francesco and Hein 2020). AutoAttack is an ensemble of diverse parameter-free attacks, including white-box and black-box attacks. In our experiments, we use standard version of AutoAttack that contains four attacks, including APGD-ce (Francesco and Hein 2020), APGD-t (Francesco and Hein 2020), FAB-t (Croce and Hein 2020), and Square (Andriushchenko et al. 2020). The perturbation budgets are $\epsilon=8$ for CIFAR-10 and Tiny-ImageNet, $\epsilon=4$ for ImageNet-1K.
4.2 Experimental Results
Comparison with SOTA Works. Table 1 compares MIMIR with previous works concerning adversarial robustness on ImageNet-1K ($l_{\infty},\epsilon=4/255$). We report the number of parameters, training epochs, steps in the inner maximization of AT, and clean and robust accuracy. The robust accuracy is evaluated by AutoAttack on the RobustBench validation set (5000 images). We divide the models into two groups: small ($\approx 22$M) and large ($\approx 86$M) models. In addition, (Singh, Croce, and Hein 2023) shows that the Convolutional stem (ConvStem) improves the robustness of ViTs. We also include an evaluation of ViTs+ConvStem in Table 2. ConvStem (Xiao et al. 2021) refers to replacing the patch embedding in ViTs with a convolutional block. MIMIR outperforms all SOTA works.
Table 2: Influence of the ConvStem block. The Standard pre-trained weights refer to pre-trained weights from timm. The configurations are deit_small_patch16_224 and vit_base_patch16_224.augreg_in1k. The weights are fine-tuned 300 epochs by 2-step APGD. †The result on ViT-S is taken from (Singh, Croce, and Hein 2023).
Architecture Pre-trained weights Natural Robust
ViT-S+ConvStem Standard 72.56 48.08†
MIMIR 72.72 48.44
ViT-B+ConvStem Standard 75.26 52.70
MIMIR 76.72 54.04
Ablation Study. Table 3 provides an ablation study to verify the design choices of MIMIR. The ablation uses 100 epochs of 1-step PGD (PGD1) AT as the baseline. Then, we apply end-to-end clean ImageNet-1K pre-training (weights available in timm library) as initialization of AT. After that, we replace the clean pre-training with MAE, adv MAE, and MIMIR step by step. The adv MAE refers to using adversarial examples but not using the MI $I(x+\delta,z)$ in loss. The pre-training scheduler is 800 epochs. We also use stronger adversarial fine-tuning for better performance, including 2-step PGD (PGD2), APGD (APGD2) FT, and a longer fine-tuning scheduler (300 epochs). Please note that a longer training scheduler does not guarantee better results (Mo et al. 2022). Our results indicate that MIMIR outperforms baselines and can be further improved under the long training scheduler.
In addition, MIMIR is also efficient and increases less than 5% of time consumption on ImageNet-1K compared to MAE, which is shown in Table 12 in the Appendix E.
@@ 4.2 Experimental Results
Architecture Training Recipe Natural Robust

### v5

@@ III-C Design Details
Autoencoder. MIMIR consists of an encoder $f_{e}$ and a decoder $f_{d}$ aligned with the general design of MAE [37]. As with other autoencoders, the encoder extracts discriminative features $z$ from inputs $x$. The decoder reconstructs the original inputs according to the discriminative features. Following the design of ViT [3], the input $x$ is separated into non-overlapping image patches. We randomly mask out a part of the patches and use the remaining patches as inputs for the following process in the encoder. This random masking process uses uniform distribution to prevent potential sampling bias, such as all foreground information being masked, as it becomes infeasible to find the reconstruction target. Thus, we aim to keep a part of the foreground as a hint for reconstruction. The information of masked content is recorded as mask tokens $m$, which are not used by the encoder but reserved for later use by the decoder. Each token is a learned vector indicating the presence of a masked patch to be predicted. The mask token is shared by all inputs of the decoder. Like unmasked patches, mask tokens are also assigned corresponding positional embeddings to be in the correct location in the reconstructed image. We emphasize that mask tokens are not used in the encoder part.
@@
To train a ViT, we use the same transformer blocks as ViT to build the encoder. The encoder only processes the visible patches, making training much more efficient. When converting to other architectures, such as ConViT [5], we use corresponding transformer blocks to build the encoder. The decoder accepts the encoded visible image patches and mask tokens as inputs. The decoder is built using the same transformer blocks as the encoder instead of using ViT [3] transformer blocks for all. Then, the decoder is followed by a fully connected layer, which outputs the same number of patches as the original image.
@@
Adversarial Pre-training Target. The training target is to extract discriminative features from visible image patches by the encoder and then reconstruct the invisible patches by the decoder. Therefore, we need a differentiable measurement to quantify the distance between the original image and the reconstructed results. Following the original MAE [37], this distance is measured by Mean Squared Error (MSE). To create a more difficult reconstruction task, we apply adversarial perturbations $\delta$ to the inputs of the encoder. Thus, the adversarial perturbations are also masked along with the image upon input into the encoder. The decoder reconstructs the original natural inputs by using the latent features $z$ extracted from adversarial examples. The outputs of decoder $x_{re}$ and $x$ are used to calculate the MSE loss, which is further used to optimize the model. Note that the reconstruction differs from the original MAE; we do not use the encoder inputs as reconstruction targets. Formally, the pre-training process (described in Algorithm 1) can be written as follows:
@@
MI as Penalty. Inspired by IB, we show in Section III-D that MI between latent representation and adversarial examples decreases as the accuracy on adversarial examples, i.e., $I(x+\delta,z)$, decreases while training. Motivated by this finding, we directly use $I(x+\delta,z)$ as a penalty in our final loss function:
@@
where $\lambda$ is a regularizer for the MI penalty. We use $I(x+\delta,z)$ instead of $I(x,z)$ as a penalty. This is because $x\to x+\delta\to z$ follows the Markov chain since $z$ is extracted from $x+\delta$. According to Data Processing Inequality (DPI) [57], $I(x,z)\leq I(x+\delta,z)$. $I(x+\delta,z)$ is closer to $z$ on the Markov chain.
Generating Adversarial Examples. To conduct the adversarial pre-training, we need an attack that finds proper adversarial perturbations $\delta$. As the autoencoder does not provide classification outputs, it is not possible to directly use existing adversarial attacks, such as PGD [21]. Nevertheless, it is feasible to design a new algorithm to find $\delta$ by maximizing $loss_{\text{mse}}$ in Eq. (7). As the feature $z$ is extracted from only visible image patches, we only attack the visible patches. We do not add any perturbations to mask tokens since the outputs of the autoencoder are only impacted by visible patches. Then, the adversarial pre-training learning objective can be written as:
@@
Figure 2: The example plots for the lower and upper bounds on the MI in Propositions III.2 and III.3. The entropy ($H(\cdot)$) is chosen uniformly at random from a set of 10 classes. The lower bound reaches its minimum at $p_{e}=0.9$.
III-D Theoretical Justification
Next, we provide theoretical justification showing that MI between the adversarial example and its latent representation, i.e., $I(x+\delta,z)$, should be constrained. Let $F$ denote any classifier trained on natural samples with desirable prediction accuracy, which may suffer from adversarial attacks. We begin our analysis by first presenting Lemma III.1.
Lemma III.1.
@@ III-D Theoretical Justification
Now, we define $p_{e}$ as the probability that the predicted label of $x+\delta$ by $F$ is not equal to that of $x_{re}$, i.e., $p_{e}=\mathbb{P}(F(x+\delta)\neq F(x_{re}))$. Intuitively, our autoencoder is trained to recover only the natural sample $x$ without any interference from $\delta$. Hence, a relatively large value of $p_{e}$ is expected. In the following, we establish the connection between $p_{e}$ and $I(x+\delta,z)$ with both lower and upper bounds, showing that minimizing $I(x+\delta,z)$ also encourages a large value of $p_{e}$.
Proposition III.2.
@@
By applying Fano’s inequality [58, 59], we obtain:
@@
The last line of Eq. (17) is by Lemma III.1. ∎
Therefore, we obtain a lower bound of $I(x+\delta,z)$. If we use CIFAR-10 ($|F(x+\delta)|=10$) and assume the predicted labels $F(x+\delta)$ follow a uniform distribution, we can visualize the lower bound as a function of $p_{e}$ as shown in Figure 2, from which we observe an obvious monotonic inverse relationship between $I(x+\delta)$ and $p_{e}$ in the range $p_{e}\in[0,0.9]$. In fact, we can also obtain an upper bound under the assumption that $I(F(x+\delta),x_{re})\approx I(x+\delta,z)$, i.e., there is no information loss in the two Markov chains in Eq. (11).
Proposition III.3.
@@
By the Hellman-Raviv inequality [60, 61], we have:
@@
Similar to the lower bound, the upper bound also indicates $I(x+\delta,z)$ is inversely proportional to $p_{e}$ as shown in Figure 2. In fact, apart from the above-mentioned lower and upper bounds, there exists an alternative and intuitive way to understand the mechanism of minimizing $I(x+\delta,z)$. For simplicity, let us assume the natural data $x$ and adversarial perturbations $\delta$ are independent [^4] This assumption is mild for certain scenarios, such as when considering universal or image-agnostic perturbations [62]., then:
@@
According to [63], minimizing the expected reconstruction error between natural sample $x$ and corrupted input $x+\delta$ amounts to maximizing a lower bound of the mutual information $I(x,z)$, even though $z$ is a function of the corrupted input. Therefore, by minimizing $I(x+\delta,z)$, the network is forced to minimize $I(\delta,z)$ (since $I(x,z)$ is maximized). In other words, only the adversarial information about $\delta$ has been removed from $z$ when minimizing $I(x+\delta,z)$. This also explains the robustness of $z$.
IV Experimental Evaluation
IV-A Experimental Setup
@@
We evaluate MIMIR on three datasets: ImageNet-1K [36], Tiny-ImageNet [35], and CIFAR-10 [34], with diverse and commonly used architectures with multiple scales: ViT [3], ConViT [5], Swin Transformer [39], and ConvNext [41]. Details of datasets are provided in Appendix -A. Hyperparameters of the decoder are included in Appendix -B.
Training Setup. We train models from scratch for all experiments. Following MAE [37], we do pre-training by MIMIR for 800 epochs. Please note that we also compare our MIMIR + fine-tuning paradigm with the End2End paradigm. The End2End paradigm refers to the supervised training of a model from scratch without self-supervised pre-training. To compare between End2End and pre-training + fine-tuning, the common training schedule practice is pre-training 800 epochs + fine-tuning 100 epochs versus End2End training 300 epochs in the existing works [37, 32, 64].
@@
The number of warmup epochs is 40 for pre-training. We use AdamW [65] as the optimizer for both pre-training and fine-tuning. We apply the cosine decay as the learning rate scheduler. At pre-training, MIMIR uses the 1-step PGD to generate adversarial examples for all three datasets. The perturbation budget is $\epsilon=8,\alpha=10$. For the fine-tuning stage, we use the 10-step PGD AT with perturbation bound $\epsilon=8,\alpha=2$ for CIFAR-10 and Tiny-ImageNet. For ImageNet-1K, the perturbation bound is $\epsilon=4$. We use different steps of PGD or APGD to generate the adversarial perturbation for training. Details on training hyperparameters are provided in Appendix -C.
TABLE I: Comparison between End2End AT and Pre-training (MIMIR) + Fine-tuning using ViT-S on CIFAR-10.
Training Method Natural PGD AutoAttack
End2End AT [21] 75.36 32.84 26.17
Fast AT [24] 76.81 32.57 21.41
TRADES [25] 74.96 32.12 24.90
MART [66] 72.42 24.47 23.45
Generalist [1] 60.88 14.44 11.20
DBAT [2] 68.32 18.83 5.25
MIMIR
@@
[…]
TABLE II: Standard deviation of 3 runs on 3 datasets.
Performance CIFAR-10 Tiny-ImageNet ImageNet-1K
[…]
TABLE III: Comparison with SOTA results on ImageNet-1K under $\epsilon=4/255$. The “Adv. Steps” refers to attack steps for generating adversarial examples for AT. The results of [31] are evaluated using 20-step PGD (AutoAttack is designed as a more powerful alternative to PGD), which is marked as $\dagger$} in the table. Although AdvXL only uses 20 epochs, it takes more time due to huge datasets for pre-training and fine-tuning.
Architecture Params (M) FT FT Epoch Adv. Steps Source Natural AutoAttack
DeiT-S 22.1 PGD 100 1 Augmentation warm-up [9] 66.62 36.56
DeiT-S 22.1 PGD 110 1 Light Recipe [20] 66.80 37.90
ViT-S 22.1 PGD 120 3 EasyRobust [67] 66.43 39.20
ViT-S 22.1 PGD 300 3 Adversarially Trained [33] 70.7 43.7
RobArch-S 26.1 PGD 110 3 RobArch [18] 70.17 44.14
ViT-S 22.1 APGD 300 2 Pre-training+AT [17] 69.22 44.04
ViT-S 22.1 PGD 300 3 MIMIR 71.52 45.90
ViT-S 22.1 APGD 300 2 MIMIR 71.00 46.10
ViT-S 22.1 APGD 300 3 MIMIR 70.96 46.16
ViT-B 86.6 ARD+PRM 10 5 ARD+PRM [11] 69.10 34.62
Swin-B 87.7 ARD+PRM 10 5 ARD+PRM [11] 74.36 38.61
ViT-B 86.6 PGD 120 3 EasyRobust [67] 70.64 43.04
Swin-B 87.7 PGD 120 3 EasyRobust [67] 75.05 47.42
RobArch-L 104 PGD 100 3 RobArch [18] 73.44 48.94
ViT-B 86.6 PGD 300 3 Adversarially Trained [33] 74.7 49.7
[…]
ViT-B 86.6 APGD 300 2 Pre-training+AT [17] 74.10 50.30
ViT-B 86.6 APGD 100 2 MIMIR 74.40 51.92
ViT-B 86.6 PGD 100 3 MIMIR 75.68 52.96
ViT-B 86.6 PGD 300 3 MIMIR 76.98 53.84
ViT-B 86.6 APGD 300 2 MIMIR 76.32 54.28
Evaluation Metrics. We use natural accuracy and robust accuracy as evaluation metrics. Natural accuracy refers to the accuracy of natural and unmodified inputs. The robust accuracy measures the accuracy under the AutoAttack (AA) [68]. AutoAttack is an ensemble of diverse parameter-free attacks, including white-box and black-box attacks. In our experiments, we use the standard version of AutoAttack that contains four attacks, including APGD-ce [68], APGD-t [68], FAB-t [69], and Square [70]. The perturbation budgets for evaluation are $\epsilon=8$ for CIFAR-10 and Tiny-ImageNet, $\epsilon=4$ for ImageNet-1K. In addition, we also evaluate the MIMIR-trained models with unforeseen attacks, such as CW attacks, attacks with $l_{2}$ norm, and out-of-distribution data (ImageNet-Corruption [42]).
Training stability. We also show that the natural accuracy and robustness of MIMIR are stable evaluation metrics. Due to the high computation cost, we cannot report the standard deviation for all experiments. To show that our method MIMIR has low variances, we train ViT-S on three datasets three times (1-step PGD AT for ImageNet-1K) and report the standard deviation and average performance in Table II.
IV-B Main Results
We first explore different AT methods and MIMIR for ViT on CIFAR-10, demonstrating the fundamental incompatibility between conventional AT approaches and ViT architectures. Following this baseline evaluation, we scale our investigation to the more challenging ImageNet-1K dataset, demonstrating the generalizability and scalability of our proposed MIMIR framework. The subsequent sections present comprehensive experimental results across three benchmark datasets: CIFAR-10, Tiny-ImageNet, and ImageNet-1K. This multi-scale evaluation strategy allows a thorough analysis of MIMIR’s effectiveness under varying conditions, from smaller to large-scale visual recognition tasks.
In addition, we also explore the effectiveness of elucidating diffusion model (EDM) data on AT in Appendix -D. This generated data is commonly used in AT to improve robustness [71, 72, 73, 74]. Specifically, we use 5 million generated CIFAR-10 data and 1 million Tiny-ImageNet data provided by [71]. Table XVIII in Appendix -D shows that EDM data significantly improves the robustness.
CIFAR-10. Table I shows the performance of End2End adversarial training from scratch and Pre-training (MIMIR) + Fine-tuning on ViT-S trained on CIFAR-10. We provide the performance of 6 established or SOTA AT methods on CIFAR-10, indicating that traditional AT training strategies are not applicable to ViTs. Importantly, our experimental results also demonstrate that MIMIR can substantially improve all AT methods. The reason is that training ViTs from scratch is known to be difficult [3, 75] and even more difficult for adversarial training [11]. For example, robust accuracy is lower than 30% on ViT-B without pre-training [11]. In contrast, MIMIR provides a more straightforward methodology and avoids this difficulty by switching to pre-training with a theoretically grounded MIM learning task.
TABLE IV: Time consumption of End2End vs. Pre-training+Fine-tuning.
Arch GPU AT Method Epochs Hours
ViT-S 4 A6000 PGD10 300 187.64
4 A6000 MIMIR(PT)+PGD10(FT) 800(PT)+100(FT) 123.76
ViT-B 4 A6000 PGD10 300 451.39
4 A6000 MIMIR(PT)+PGD10(FT) 800(PT)+100(FT) 263.77
Swin-L 4 H100 PGD3 300 363.42
4 H100 MIMIR(PT)+PGD3(FT) 800(PT)+100(FT) 231.14
TABLE V: The performance of MIMIR on ImageNet-1K against unforeseen threats. The $l_{2}$ version of the CW [76] attack is limited by $c$. Higher $c$ allows a more powerful perturbation.
Architecture Method Natural CW ($l_{2},c=1$) CW ($l_{\infty},\epsilon=4/255$) AA ($l_{2},\epsilon=2.0$) AA ($l_{\infty},\epsilon=4/255$) ImageNet-C [42]
ViT-S [17] 69.22 44.28 45.98 37.52 44.04 45.62
MIMIR 70.96 66.34 49.22 49.12 46.16 48.78
ViT-B [17] 74.10 59.42 52.76 52.12 50.30 53.69
MIMIR 76.32 65.30 56.82 56.72 54.28 57.07
Time Consumption (End2End vs. Pre-training(PT)+Fine-tuning(FT)). Note that we follow the standard way to compare End2End and MIMIR (Pre-training + Fine-tuning) training methods by fixing the training schedule, following existing works [37, 32, 64, 77, 40], where we include pre-training 800 epochs + fine-tuning 100 epochs versus supervised End2End training of 300 epochs. The reason for having a larger number of pre-training epochs is that self-supervised pre-training is much more efficient (see Table XXII in Appendix -I) than End2End training, and the pre-trained backbone can be used multiple times for various fine-tuning tasks. For example, in Table I, the 6 End2End AT methods cost $300\times 6=1800$ fine-tuning epochs. MIMIR costs 800 pre-training epochs + $100\times 6$ fine-tuning epochs, i.e., we only conduct the pre-training once for results in Table I. MIMIR pre-training epoch is more efficient than a fine-tuning epoch by discarding 75% of image patches.
More specifically, Table IV shows the total time consumption of End2End vs. Pre-training+Fine-tuning on three architectures with ImageNet-1K. Although MIMIR takes more training epochs, its pre-training+fine-tuning paradigm still costs less time than End2End adversarial training and gains much better performance on both natural and adversarial examples. Additional time consumption results with different datasets can be found in Table XXII in Appendix -I.
ImageNet-1K. Table III compares MIMIR with previous works concerning adversarial robustness on ImageNet-1K ($l_{\infty},\epsilon=4/255$), which follows the evaluation of common standardized RobustBench [78]. Similar to other works [20, 33, 17], we consider simpler AT methods (i.e., PGD and APGD AT) instead of the latest AT methods, such as Generalist [1] and DBAT [2]. Indeed, since the latest methods introduce tailored components for CNNs to improve their adversarial robustness, they might not be effective for ViTs. The number of parameters, training epochs, steps in the inner maximization of AT, and clean and robust accuracy are reported to provide a more detailed understanding of the performance. The robust accuracy is evaluated by AutoAttack on the RobustBench [78] validation set (5,000 images). We divide the models into: small ($\approx 22$M) and large ($\approx 86$M) models, corresponding to ViT-S and ViT-B. Experimental results demonstrate that MIMIR outperforms all previous works across various training setups.
MIMIR with Various Architectures. In Table VI, we show that MIMIR can be applied to diverse architectures. Specifically, we use three representative options, including ViT+convolutional blocks (CVST) [17], the latest CNN architecture (ConvNext [41]), and a hierarchical vision transformer (Swin [39]). The ViT+CVST refers to using ConvStem [79] to replace the patch embedding in ViTs with a convolutional block. The ViT+CVST shows improved robustness compared to pure ViT models according to experiments in [17].
As CNN and hierarchical architecture cannot accept variable-length inputs, MIMIR is not directly compatible with ConvNext and Swin. To adapt MIMIR to the hierarchical Swin Transformer, we implemented Masked Image Modeling using Group Window Attention [38], which groups image patches within each local window of arbitrary size and performs masked self-attention in each group. To apply MIMIR to ConvNext, we use SparK [40] for CNN to handle irregular and randomly masked input images, which is achieved by sparse convolution. MIMIR achieves better or comparable results compared to SOTA results on RobustBench [78].
MIMIR against Unforeseen Attacks. Except for adversaries who are limited by the adversarial budget, e.g., $l_{\infty}=8\text{ or }4$ or from PGD-family in our main experiments, MIMIR also shows the potential to provide robustness against unforeseen attacks and naturally corrupted data (ImageNet-C [42]). Table V demonstrates the robustness of MIMIR against practical unforeseen attacks, including non-PGD attacks ($l_{2}$ and $l_{\infty}$ CW attack [76]), $l_{2}$ AutoAttack, and ImageNet-C [42]. It is clear that MIMIR still performs well against these unforeseen attacks. In particular, MIMIR achieved top accuracy compared to the results of the ImageNet-C Leaderboard on the RobustBench.
IV-C Ablation Study and Further Analysis
Step by step ablation. Table VII provides an ablation study to verify the design choices of MIMIR. The ablation uses 100 epochs of 1-step PGD (PGD1) AT as the baseline. Then, we apply end-to-end clean ImageNet-1K pre-training (weights available in timm library [^5] https://github.com/huggingface/pytorch-image-models/blob/main/timm/models/vision_transformer.py) as initialization of AT. After that, we replace the clean pre-training with MAE, adv MAE, and MIMIR step by step. The adv MAE refers to using adversarial examples but not using the MI $I(x+\delta,z)$ in the loss. The pre-training schedule is 800 epochs. We also use stronger adversarial fine-tuning for better performance, including 2-step PGD (PGD2), APGD (APGD2) FT, and a longer fine-tuning scheduler (300 epochs). Our results indicate that MIMIR outperforms baselines and can be further improved under the long training schedule.
Figure 3: Natural, adversarial accuracy, and MIMIR pre-training loss of ViT-S with different numbers of pre-training epochs under a 20-step PGD attack. The performance increases as the number of pre-training epochs increases (loss decreases).
TABLE VI: Comparsion with SOTA ImageNet-1K results on RobustBench [78] with different architectures. $\dagger$}: The CVST modules are also pre-trained with MIMIR.
Architecture Method FT Epoch Natural AutoAttack
ViT-S+CVST [17] 300 72.56 48.08
MIMIR 300 72.72 48.44
[…]
ViT-B+CVST [17] 250 76.30 54.66
MIMIR 300 76.72 54.04
[…]
ConvNext-T [17] 300 72.40 48.60
MIMIR 300 72.50 48.76
Swin-B [33] 300 76.16 56.16
MIMIR 150 76.62 55.90
Swin-L [33] 300 78.92 59.56
MIMIR 100 78.62 59.68
Longer epochs (lower loss) provide better performance. Pre-training epoch. Our experimental framework employs an 800-epoch pre-training as the baseline configuration. To systematically evaluate the impact of training duration, we conduct a comprehensive ablation study using ViT-S architectures, varying the number of pre-training epochs while maintaining a fixed 50-epoch fine-tuning for all models. As demonstrated in Figure 3, we observe several key phenomena. Extended pre-training schedules consistently yield lower MIMIR loss values, and this reduction in loss results in measurable improvements in both adversarial and natural accuracy. In addition, the improved performance with loss MIMIR loss also indicates that the model capacity is not saturated within the tested epoch range and could be further improved with a larger number of epochs.
Fine-tuning epoch. To isolate and quantify the contribution of MIMIR pre-training to model performance, we employ a short fine-tuning for the pre-trained models. This is to train the randomly initialized classification layer since we do not have the classification layer at pre-training. This approach allows us to evaluate the quality of the representations learned during pre-training. In Figure 5, we show that MIMIR pre-training plus 5 or 10 epochs of fine-tuning is enough to achieve similar performance compared to 100-epoch fine-tuning. These results suggest that the majority of the model’s final performance is attributable to the MIMIR pre-training phase.
TABLE VII: Ablation of pre-training (PT) and fine-tuning (FT) methods on ImageNet-1K. ($\dagger$}: catastrophic over-fitting [24] due to 1-step AT when fine-tuning, which can be fixed by 2-step AT. The fixed natural and robust accuracy are 69.96 and 36.90, respectively.)
Architecture Training Recipe Natural AutoAttack